# ST-Booster: An Iterative SpatioTemporal Perception Booster for Vision-and-Language Navigation in Continuous Environments

### v1

@@ III-C Multi-Granularity Aligned Fusion Module
Although global topological and local grid maps each possess distinct perceptual advantages, their structural limitations still constrain their perception capability. Therefore, this section establishes a multi-granularity alignment fusion module, which introduces instruction-aware feature alignment to construct unified hybrid representations aligned with linguistic instructions. Meanwhile, the cross-domain fusion between long- and short-term perception domains addresses intrinsic heterogeneity in spatiotemporal feature spaces.
Instruction Feature Alignment. By interacting with instruction information, contextual understanding is embedded into domain-specific environmental perception. We adopt Topology-based Cross-Modal Transformer (TCMT) and Grid-based Cross-Modal Transformer (GCMT)  [18] to facilitate the fusion of instruction cues with long-term and short-term environmental perception, respectively. Specifically, TCMT incorporates a graph-aware self-attention mechanism to integrate instructions $\mathbf{I}$ with global topological information $\mathbf{{G}}_{t}$, producing a global-aware multi-modal graph $\mathbf{\tilde{G}}_{t}$, i.e.,
$\mathbf{\tilde{G}}_{t}=\text{TCMT}(\text{NE}(\mathbf{{G}}_{t}),\mathbf{I}),$ (1)
where NE()($\cdot$)( ⋅ ) represents the node embedding network. For GCMT, the fusion of instruction information $\mathbf{I}$ and local spatial awareness $\mathbf{M}_{t}$ is achieved through a self-attention mechanism for cell encoding, i.e.,
$\mathbf{\tilde{M}}_{t}=\text{GCMT}(\text{CE}(\mathbf{M}_{t}),\mathbf{I}),$ (2)
where $\mathbf{\tilde{M}}_{t}$ represents the local-aware multi-modal grid maps and CE()($\cdot$)( ⋅ ) represents the grid cell embedding network.
@@
Cross-Domain Perception Fusion. To leverage the complementary strengths of both map structures, we introduce a geometrically-constrained projection operator to facilitate cross-domain knowledge sharing. The operator comprises two components, Cell2Node and Node2Cell, which project perceptual information from heterogeneous domains into a unified feature space. As depicted in the bottom-right corner of Figure 3, Cell2Node is initially utilized to transfer local grid features into the long-term topological domain. To maintain a domain-invariant structure, suitable feature migration strategies are applied based on the different node types. In particular, grid maps are aggregated to provide comprehensive environmental information for the visited nodes $\mathbf{n}$. As for the unvisited nodes $\mathring{\mathbf{n}}$ with partly observations, we locate them on the grid cells and extract the corresponding features. With the utilization of Cell2Node, the intermediate topological map $\mathbf{G}^{\prime}_{t}$ is generated, with the visited nodes $\mathbf{n}^{\prime}$ and observed nodes $\mathring{\mathbf{n}}^{\prime}$ with new representations from $\mathbf{M}_{t}$ as follows
$\mathbf{n}_{i}^{\prime}=\text{average}(\mathbf{M}_{i}),i=1,\cdots,t,$ (3)
$\mathring{\mathbf{n}^{\prime}}^{j}_{i}=\mathbf{m}_{x_{i}^{j},y_{i}^{j}},i=1,\cdots,t;j=1,\cdots,g_{i},$ (4)
@@
where average()($\cdot$)( ⋅ ) represents an expectation pooling operation, $x_{i}^{j}$ and $y_{i}^{j}$ denote the geometric coordinates of the $j-$th observed node at time step $i$ when projected onto the grid map, $g_{i}$ is the observed candidate number at $i$ step and $\mathbf{m}_{x_{i}^{j},y_{i}^{j}}$ represents the grid cell features with the coordinate $(x_{i}^{j},y_{i}^{j})$ in grid map $\mathbf{M}_{i}$. Subsequently, $\mathbf{G}^{\prime}_{t}$ is fused with the original graph $\mathbf{G}_{t}$ to form a hybrid graph, i.e.,
$\hat{\mathbf{G}}_{t}=\text{GF}([\mathbf{G}^{\prime}_{t};\mathbf{G}_{t}]),$ (5)
@@
where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes the concatenation operation, and GF()($\cdot$)( ⋅ ) is a fully-connected network fusing the integrated graph features. Subsequently, $\hat{\mathbf{G}}_{t}$ is utilized to replace the original topological map $\mathbf{G}_{t}$ in Equation 1 with more fine-grained spatial information.
@@
Similarly, we introduce Node2Cell to project the global historical perception of nodes into the local grid-based domain, which is shown in the upper right corner of Figure 3. To effectively capture comprehensive global context, such as task progress, we leverage linguistically-aligned topological maps $\tilde{\mathbf{G}}_{t}$ with nodes $\tilde{\mathbf{n}}_{t}$ in Equation 1 to facilitate cross-domain knowledge fusion. Specifically, we flatten the node features and map them back into the grid maps with same geometric positions. To reduce perceptual bias from isolated node features, we fuse multiple nodes within the grid map’s spatial boundaries, represented as $\mathcal{N}_{t}=\{\tilde{\mathbf{n}}_{i}|\text{dis}(\tilde{\mathbf{n}}_{t},\tilde{\mathbf{n}}_{i})\text{\textless}U_{b}\},\tilde{\mathbf{n}}_{i}\in\mathcal{V}^{\text{v}}_{t}$, where dis(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) calculates the Euclidean distance between two nodes, and $U_{b}$ represents the geometry size of the grid maps. We also introduce a geometric-aware discount matrix $\mathbf{D}_{i}$ to adjust the contribution of multiple projected node features to each grid cell. Specifically, the nodes in $\mathcal{N}_{t}$ are projected onto the current grid map with coordinates $(x_{n},y_{n}),n=1,\dots,\text{len}(\mathcal{N}_{t})$. The Euclidean distances $d_{n}^{j},j=0,\dots,U\times U-1$ from the $j-$th grid cell to $(x_{n},y_{n})$ are calculated, where the maximum and minimum distance are represented as $d_{n}^{\text{max}}$ and $d_{n}^{\text{min}}$, respectively. Through the distance normalization, the discount matrix $\mathbf{D}_{i}$ is obtained by $\mathbf{D}_{i}=\{\frac{d_{n}^{\text{max}}-d_{n}^{j}}{d_{n}^{\text{max}}-d_{n}^{\text{min}}}\}$. The intermediate perception integrated from the grid map is further updated from the multi-modal grid maps $\tilde{\mathbf{M}}_{t}$ obtained in Equation 2, i.e.,
$\hat{\mathbf{M}}_{t}=\text{MF}([(\sum_{i}\mathbf{D}_{i}\times\text{flatten}(\tilde{\mathbf{n}}_{i}));\tilde{\mathbf{M}}_{t}]),$ (6)
@@
where MF()($\cdot$)( ⋅ ) serves as a fusion network of grid map features. Through bidirectional geometric-aware transformation, topological graphs enriched with geometric details, while grid maps enhanced with global context, and the multi-granularity map representations are constructed.
@@ III-D Value-Guided Waypoint Generation Module
We introduce the guided attention heatmap (GAH) prediction task to improve model interpretation by explicitly capturing the alignment between hybrid map perception and instructions. GAH is a self-centered heatmap that predicts the 2D probability distribution of instruction-relevant regions, derived from multi-granularity grid map $\mathbf{\tilde{M}}_{t}$ representations, i.e.,
$\mathbf{H}_{t}=\textrm{HFFN}(\mathbf{\tilde{M}}_{t}),$ (7)
where HFFN()($\cdot$)( ⋅ ) is a fully-connected forward network to predict GAHs. To obtain the ground truth heatmap $\mathbf{H}^{*}_{t}$, we approximate the relevant region based on the position of the next waypoint $\mathbf{w}^{*}_{t+1}$ in the reference trajectory, without additional semantic information. Specifically, we project $\mathbf{w}^{*}_{t+1}$ onto an egocentric coordinate system aligned with agent’s current heading, and smooth the located position with the Gaussian kernel. To accurately capture the relationship between the waypoint and the surrounding environment, the map’s resolution is enhanced by $m\times$ through grid upsampling, resulting in the $(mU)\times(mU)$ higher-definition grid map. Based on the ground-truth GAHs, the weakly-supervised training method discussed in Section III-E is applied to improve the module to capture contextual cues by leveraging multi-modal cross-domain knowledge from hybrid environmental perception.
To overcome the common issue where the standard waypoint predictor [21] ignores connections between instructions and environments, we introduce GAHs to generate candidate waypoints with crucial spatial guidance considering environmental perception and language instructions. As shown in Algorithm 1, the initial waypoint distributions $\mathbf{P}_{t}$ are predicted by the original predictor WP()($\cdot$)( ⋅ ) with only visual observations, and then refined with predicted GAHs $\mathbf{H}_{t}$ in Equation 7 through weighted fusion. $\mathbf{H}_{t}$ automatically adjust the distribution heatmaps and generate candidate waypoints $\hat{\mathbf{w}}^{t}_{k}$ to the area relevant to instruction with high probability. This approach improves environmental awareness of critical zones to support downstream navigation planning, balancing knowledge from both perceptual reality and task context.
Input: Pre-trained ‘HFFN()($\cdot$)( ⋅ )’; fine-tuned VLN-CE agent ‘ST-Booster’ with detail in Section III-C; waypoint predictor ‘WP()($\cdot$)( ⋅ )’; vision encoder ‘Encoder()($\cdot$)( ⋅ )’ [39],  [40]; map update operation ‘Update()($\cdot$)( ⋅ ) with new observations, illustrated in Section III-B’; hybrid action predictor ‘AP()($\cdot$)( ⋅ )’ in Section III-E; add new observation to history path operation Add()($\cdot$)( ⋅ ); environmental operation Execute()($\cdot$)( ⋅ ); instruction embeddings $\mathbf{I}$; initial Observation $\mathbf{O}_{0}$; initial path $\mathbf{P}_{0}={\mathbf{O}_{0}}$; initial maps $\mathbf{G}_{0}$ and $\mathbf{M}_{0}$; max time step $t_{m}$ and GAH weight $\delta$.
@@
for $t=1$ to $T\leq t_{m}$  do
      $\mathbf{V_{t}}$ = Encoder($\mathbf{O}_{t}$);
      $\mathbf{P}_{t}=\text{WP}(\mathbf{V_{t}})$;
      $\mathbf{w}^{t}_{k}=\text{Sample}(\mathbf{P}_{t}),k=1,...,g_{t}$; % Sample original candidate waypoints
       $\mathbf{G}^{\text{ori}}_{t},\mathbf{M}^{\text{ori}}_{t}=\text{Update}(\mathbf{G}_{t-1},\mathbf{M}_{t-1},\mathbf{w}^{t}_{k})$;
       $\hat{\mathbf{G}}^{\text{ori}}_{t},\hat{\mathbf{M}}^{\text{ori}}_{t}=\text{ST-Booster}(\mathbf{G}^{\text{ori}}_{t},\mathbf{M}^{\text{ori}}_{t},\mathbf{I})$;
       $\mathbf{H}_{t}=\text{HFFN}(\hat{\mathbf{M}}_{t})$; % Generate GAHs
       $\hat{\mathbf{H}}_{t}=\delta\cdot\mathbf{H}_{t}+\mathbf{P}_{t}$; % Weighted fusion
       $\hat{\mathbf{w}}^{t}_{k}=\text{Sample}(\hat{\mathbf{H}}_{t}),k=1,...,g^{\prime}_{t}$; % Sample adjusted candidate waypoints
       $\mathbf{G}_{t},\mathbf{M}_{t}=\text{Update}(\hat{\mathbf{G}}^{\text{ori}}_{t},\hat{\mathbf{M}}^{\text{ori}}_{t},\hat{\mathbf{w}}^{t}_{k})$;
       $\hat{\mathbf{G}}_{t},\hat{\mathbf{M}}_{t}=\text{ST-Booster}(\mathbf{G}_{t},\mathbf{M}_{t},\mathbf{I})$; % Choose next target waypoint.
       $\mathbf{a}_{t}=\text{AP}(\hat{\mathbf{G}}_{t},\hat{\mathbf{M}}_{t})$;
      if $\mathbf{a}_{t}\rightarrow\text{stop}$ then
            Break;
       end if
      $\mathbf{O}_{t+1}$ = Execute($\mathbf{a}_{t}$); % Execute action in the environment
      $\mathbf{P}_{t+1}=\text{Add}(\mathbf{P}_{t},\mathbf{O}_{t+1})$
end for

### v2

@@ III-C Multi-Granularity Aligned Fusion Module
Although global topological and local grid maps offer complementary perceptual strengths, their structural limitations still hinder unified scene understanding. To overcome this, we adopt an asymmetric fusion strategy within a multi-granularity aligned fusion module that performs instruction feature alignment and cross-domain perception fusion to bridge heterogeneous long- and short-term perceptions. In this design, topological features are first enriched with fine-grained grid information before instruction alignment, whereas grid features are aligned with instructions first and then fused with globally aligned task cues to preserve local geometric detail while incorporating global semantics.
Instruction Feature Alignment. By interacting with instruction information, contextual understanding is embedded into domain-specific environmental perception. We adopt Topology-based Cross-Modal Transformer (TCMT) and Grid-based Cross-Modal Transformer (GCMT)  [an2023bevbert] to facilitate the fusion of instruction cues with long-term and short-term environmental perception, respectively. Specifically, TCMT incorporates a graph-aware self-attention mechanism to integrate instructions $\mathbf{I}$ with the global topological representation enhanced by grid features ($\mathbf{G}^{\text{f}}_{t}$ introduced in Equation 5), yielding a globally informed multi-modal graph $\tilde{\mathbf{G}}^{\text{f}}_{t}$, i.e.,
$\tilde{\mathbf{G}}^{\text{f}}_{t}=\text{TCMT}(\text{NE}(\mathbf{{G}}^{\text{f}}_{t}),\mathbf{I}),$ (1)
where NE()($\cdot$) represents the node embedding network. For GCMT, the fusion of instruction information $\mathbf{I}$ and local spatial awareness $\mathbf{M}_{t}$ is achieved through a self-attention mechanism for cell encoding, i.e.,
$\tilde{\mathbf{M}}_{t}=\text{GCMT}(\text{CE}(\mathbf{M}_{t}),\mathbf{I}),$ (2)
where $\tilde{\mathbf{M}}_{t}$ represents the local-aware multi-modal grid maps and CE()($\cdot$) represents the grid cell embedding network.
@@
Cross-Domain Perception Fusion. To leverage the complementary strengths of both map structures, we introduce a geometrically-constrained projection operator to facilitate cross-domain knowledge sharing. The operator comprises two components, Cell2Node and Node2Cell, which project perceptual information from heterogeneous domains into a unified feature space. As depicted in the upper-right corner of LABEL:fig:feat, Cell2Node is initially utilized to transfer local grid features into the long-term topological domain. To maintain a domain-invariant structure, suitable feature migration strategies are applied based on the different node types. In particular, grid maps are aggregated to provide comprehensive environmental information for the visited nodes $\mathbf{n}$. As for the unvisited nodes $\mathring{\mathbf{n}}$ with partly observations, we locate them on the grid cells and extract the corresponding features. With the utilization of Cell2Node, the intermediate topological map $\mathbf{G}^{\prime}_{t}$ is generated, consisting of the visited nodes $\mathbf{n}^{\prime}$ and observed nodes $\mathring{\mathbf{n}}^{\prime}$ with new representations from $\mathbf{M}_{t}$ as follows,
$\mathbf{n}_{t}^{\prime}=\text{average}(\mathbf{M}_{t}),$ (3)
$\mathring{\mathbf{n}}^{\prime}_{t,i}=\mathbf{m}_{(u_{t,i},v_{t,i})},i=1,\cdots,g_{t},$ (4)
@@
where average()($\cdot$) represents an expectation pooling operation, $u_{t,i}$ and $v_{t,i}$ denote the geometric coordinates of the $i$-th observed node at time step $t$ when projected onto the grid map, $g_{t}$ is the observed candidate number and $\mathbf{m}_{(u_{t,i},v_{t,i})}$ represents the grid cell features with the coordinate $(u_{t,i},v_{t,i})$ in grid map $\mathbf{M}_{t}$. Subsequently, $\mathbf{G}^{\prime}_{t}$ is fused with the original graph $\mathbf{G}_{t}$ to form a hybrid graph with complete perception, i.e.,
$\mathbf{G}^{\text{f}}_{t}=\text{GF}([\mathbf{G}^{\prime}_{t};\mathbf{G}_{t}]),$ (5)
@@
where $\cdot$$\cdot$] denotes the concatenation operation, and GF()($\cdot$) is a fully-connected network fusing the integrated graph features. Subsequently, the fine-grained enhanced $\mathbf{G}^{\text{f}}_{t}$ is aligned with the instruction in Equation 1 to obtain $\tilde{\mathbf{G}}^{\text{f}}_{t}$.
@@
Similarly, we introduce Node2Cell to project the historical perception of nodes into the local grid-based domain, which is illustrated in the bottom-right corner of LABEL:fig:feat. To effectively capture comprehensive global context, such as task progress, we leverage the enhanced linguistically-aligned topological maps $\tilde{\mathbf{G}}^{\text{f}}_{t}$. Their node features $\tilde{\mathbf{n}}^{\text{f}}_{t}$ are sequentially processed by Equation 5 and Equation 1 to facilitate cross-domain knowledge fusion. Specifically, we flatten the node features and map them back into the grid maps with the same geometric positions. To reduce perceptual bias from isolated node features, we fuse multiple nodes within the grid map’s spatial boundaries, represented as $\mathcal{N}_{t}=\{\tilde{\mathbf{n}}^{\text{f}}_{t,i}|\text{dis}(\tilde{\mathbf{n}}^{\text{f}}_{t},\tilde{\mathbf{n}}^{\text{f}}_{t,i})\text{\textless}U_{b}\},\tilde{\mathbf{n}}^{\text{f}}_{t,i}\in\mathcal{V}^{\text{v}}_{t}$, where dis(,)($\cdot$$\cdot$) calculates the Euclidean distance between two nodes, and $U_{b}$ represents the geometry size of the grid maps. We also introduce a geometric-aware discount matrix $\mathbf{D}_{t,i}$ to adjust the contribution of multiple projected node features to each grid cell. Specifically, the nodes in $\mathcal{N}_{t}$ are projected onto the current grid map with coordinates $(u_{t,i},v_{t,i}),i=1,\dots,\text{len}(\mathcal{N}_{t})$. We calculate the Euclidean distances $d_{t,i,n}$ $(n=0,\dots,U\times V-1)$ between each grid cell and the grid at the map center, where $d_{t,i}^{\text{min}}$ and $d_{t,i}^{\text{max}}$ denote the minimum and maximum distances, respectively. Through the distance normalization, the discount matrix $\mathbf{D}_{t,i}$ is obtained by $\mathbf{D}_{t,i}=\{\frac{d_{t,i}^{\text{max}}-d_{t,i,n}}{d_{t,i}^{\text{max}}-d_{t,i}^{\text{min}}}\}$. The intermediate perception is further updated from the multi-modal grid maps $\tilde{\mathbf{M}}_{t}$ obtained in Equation 2, i.e.,
$\tilde{\mathbf{M}}^{\text{f}}_{t}=\text{MF}([(\sum_{i}\mathbf{D}_{t,i}\times\text{flatten}(\tilde{\mathbf{n}}^{\text{f}}_{t,i}));\tilde{\mathbf{M}}_{t}]),$ (6)
@@
where MF()($\cdot$) serves as a fusion network of grid map features. Through bidirectional geometric-aware transformation, topological graphs are enriched with geometric details, grid maps are enhanced with global context, and multi-granularity map representations are constructed.
@@ III-D Value-Guided Waypoint Generation Module
We introduce the Guided Attention Heatmap (GAH) prediction task to enhance model interpretability by explicitly aligning hybrid map perception with navigation instructions. GAH is a self-centered heatmap that predicts the 2D probability distribution of instruction-relevant regions from multi-granularity grid map representations $\tilde{\mathbf{M}}^{\text{f}}_{t}$,
$\mathbf{H}_{t}=\textrm{HFFN}(\tilde{\mathbf{M}}^{\text{f}}_{t}),$ (7)
where HFFN()($\cdot$) denotes a fully connected forward network. The ground-truth heatmap is derived from the next target waypoint in the reference trajectory, projected onto an egocentric coordinate system aligned with the agent’s heading and smoothed using a Gaussian kernel. To better capture detailed spatial relationships, the map resolution is upsampled by $m\times n$, resulting in a $(mU)\times(nV)$ high-definition grid. Based on these ground-truth GAHs, a weakly supervised training scheme (see Section III-E) is used to help the model learn contextual cues from hybrid multimodal perception.
We integrate GAH as a plug-and-play module in the evaluation phase. As illustrated in Algorithm 1, ST-Booster consists of two stages: prior guidance and fusion-based decision making. In the first stage, the waypoint predictor WP()($\cdot$) generates an initial distribution $\mathbf{P}_{t}$, which is refined through weighted fusion with the predicted $\mathbf{H}_{t}$ from Eq. (7). Superscripts in Algorithm 1 indicate features of different stages (e.g., ‘s1’ and ‘s2’ denote the first and second stages, respectively) but are omitted here for clarity. $\mathbf{H}_{t}$ adaptively adjusts the heatmap weights, guiding candidate waypoints $\hat{\mathbf{w}}_{t,k}$ toward instruction-relevant regions. In the second stage, the model performs joint global–local reasoning based on the refined candidate distribution and fused features. It maintains long-term consistency through global topological cues while optimizing spatial precision using local grids, and finally selects the target waypoint from both current and historical candidates.
III-E Pretraining and Fine-Tuning
Pretraining. To enhance the representational capability of the hybrid multi-modal maps, we introduce three pretraining tasks.
1
Input: Pre-trained ‘HFFN()($\cdot$)’ in Section III-D; fine-tuned agent ‘ST-Booster’ in Section III-C; waypoint predictor ‘WP()($\cdot$)’; vision encoder ‘Encoder()($\cdot$)’ [dosovitskiy2020image],  [he2016deep]; map update operation ‘Update()($\cdot$)’ in Section III-B; hybrid action predictor ‘AP()($\cdot$)’ in Equation 9 and Equation 10 of Section III-E; add new observation to history path operation ‘Add()($\cdot$)’; environmental operation ‘Execute()($\cdot$)’; instruction embeddings $\mathbf{I}$; initial Observation $\mathbf{O}_{0}$; initial path $\mathbf{P}_{0}={\mathbf{O}_{0}}$; initial maps $\mathbf{G}_{0}$ and $\mathbf{M}_{0}$; max time step $t_{m}$ and GAH weight $\delta$.
@@ III-E Pretraining and Fine-Tuning
2
3for $t=1$ to $T\leq t_{m}$ do
4
5 $\mathbf{V_{t}}$ = Encoder($\mathbf{O}_{t}$);
6 $\mathbf{P}_{t}=\text{WP}(\mathbf{V_{t}})$;
7 $\mathbf{w}_{t,k}=\text{Sample}(\mathbf{P}_{t}),k=1,...,g_{t}$; % Sample original candidate waypoints
8 $\mathbf{G}^{\text{s1}}_{t},\mathbf{M}^{\text{s1}}_{t}=\text{Update}(\mathbf{G}_{t-1},\mathbf{M}_{t-1},\mathbf{w}_{t,k})$;
9 $\tilde{\mathbf{G}}^{\text{f-s1}}_{t},\tilde{\mathbf{M}}^{\text{f-s1}}_{t}=\text{ST-Booster}(\mathbf{G}^{\text{s1}}_{t},\mathbf{M}^{\text{s1}}_{t},\mathbf{I})$;
10 $\mathbf{H}_{t}=\text{HFFN}(\tilde{\mathbf{M}}^{\text{f-s1}}_{t})$; % Generate GAHs
11 $\hat{\mathbf{H}}_{t}=\delta\cdot\mathbf{H}_{t}+\mathbf{P}_{t}$; % Weighted fusion
12 $\hat{\mathbf{w}}_{t,k}=\text{Sample}(\hat{\mathbf{H}}_{t}),k=1,...,g^{\prime}_{t}$; % Sample adjusted candidate waypoints
13 $\mathbf{G}^{\text{s2}}_{t},\mathbf{M}^{\text{s2}}_{t}=\text{Update}(\mathbf{G}_{t-1},\mathbf{M}_{t-1},\hat{\mathbf{w}}_{t,k})$;
14 $\tilde{\mathbf{G}}^{\text{f-s2}}_{t},\tilde{\mathbf{M}}^{\text{f-s2}}_{t}=\text{ST-Booster}(\mathbf{G}^{\text{s2}}_{t},\mathbf{M}^{\text{s2}}_{t},\mathbf{I})$;
15 $\mathbf{a}_{t}=\text{AP}(\tilde{\mathbf{G}}^{\text{f-s2}}_{t},\tilde{\mathbf{M}}^{\text{f-s2}}_{t})$;
16 if $\mathbf{a}_{t}\rightarrow\text{stop}$ then
17    Break;
18
19 $\mathbf{O}_{t+1}$ = Execute($\mathbf{a}_{t}$); % Execute action in the environment
20 $\mathbf{P}_{t+1}=\text{Add}(\mathbf{P}_{t},\mathbf{O}_{t+1})$